OpenStudy (anonymous):

A force of 110N is applied horizontally to the handles of a lawnmower to move it at a speed of 0.80m/s across a lawn, Find the power used to mow the lawn.

sam (.sam.):

P=Fv

sam (.sam.):

P=110(0.8)=88W

OpenStudy (anonymous):

since the force and the displacement are in the same direction, the work done is positive. Also Power=work/time. So P=f*d/t P=f*v P=110*0.8= 88 watt

OpenStudy (anonymous):

a=v-u/t 10=0-0.8/t t=0.8/10 t=0.08 p=workdone/time p=110/0.08 p=110/0.08